# A Scalable Unsupervised Framework for multi-aspect labeling of Multilingual and Multi-Domain Review Data

### v1

@@ Appendix D Prompt
Figure D.1: Zero-shot Prompt for Hotel Aspect Category Labeling (Korean with English Translation).This figure shows a zero-shot prompt used by the LLM to categorize hotel-related review sentences into predefined aspects (e.g., purpose, room, location, facilities, satisfaction, view, service). The original instructions and labeling guidelines are in Korean, with an English translation provided for clarity.
@@
Figure D.2: Zero-shot Food Prompt (Korean with English Translation). This figure presents a zero-shot prompt used by the LLM to categorize food-related review sentences into predefined aspects (e.g., satisfaction, taste, purpose, atmosphere, service, location, food, portion size, waiting time). The instructions and labeling guidelines are originally written in Korean, with an English translation included for clarity.
@@
Figure D.3: Zero-shot Beauty Prompt (English Only). This figure shows a zero-shot prompt used by the LLM to classify beauty-related review sentences into predefined aspects (e.g., color, hair, improvement, packaging, etc.). All instructions, aspect lists, and labeling guidelines are provided in English, and users must select only from the given aspect categories without adding any unlisted keywords.
@@
Figure D.4: Few-Shot Examples for the Hotel & Food Domains. This figure demonstrates how examples of multi-aspect category labels assigned to reviews in the Hotel and Food domains are provided to the model in a few-shot setting. By referring to the aspect categories assigned to each review, the model is guided to classify new sentences more accurately.
@@
Figure D.5: Few-Shot Examples for the Beauty Domains. This figure demonstrates how examples of multi-aspect category labels assigned to reviews in the Beauty domains are provided to the model in a few-shot setting. By referring to the aspect categories assigned to each review, the model is guided to classify new sentences more accurately.

### v2

@@ Appendix E Prompt
Figure E.1: Zero-shot Prompt for Hotel Aspect Category Labeling (Korean with English Translation).This figure shows a zero-shot prompt used by the LLM to categorize hotel-related review sentences into predefined aspects (e.g., purpose, room, location, facilities, satisfaction, view, service). The original instructions and labeling guidelines are in Korean, with an English translation provided for clarity.
@@
Figure E.2: Zero-shot Food Prompt (Korean with English Translation). This figure presents a zero-shot prompt used by the LLM to categorize food-related review sentences into predefined aspects (e.g., satisfaction, taste, purpose, atmosphere, service, location, food, portion size, waiting time). The instructions and labeling guidelines are originally written in Korean, with an English translation included for clarity.
@@
Figure E.3: Zero-shot Beauty Prompt (English Only). This figure shows a zero-shot prompt used by the LLM to classify beauty-related review sentences into predefined aspects (e.g., color, hair, improvement, packaging, etc.). All instructions, aspect lists, and labeling guidelines are provided in English, and users must select only from the given aspect categories without adding any unlisted keywords.
@@
Figure E.4: Few-Shot Examples for the Hotel & Food Domains. This figure demonstrates how examples of multi-aspect category labels assigned to reviews in the Hotel and Food domains are provided to the model in a few-shot setting. By referring to the aspect categories assigned to each review, the model is guided to classify new sentences more accurately.
@@
Figure E.5: Few-Shot Examples for the Beauty Domains. This figure demonstrates how examples of multi-aspect category labels assigned to reviews in the Beauty domains are provided to the model in a few-shot setting. By referring to the aspect categories assigned to each review, the model is guided to classify new sentences more accurately.